OpenStudy (anonymous):

derivative sqrt((3-x)(2+x))+5 arcsin*sqrt((x+2)/5) Please help :)

OpenStudy (anonymous):

\[\sqrt{(3-x)(2+x)}+5\sin^{-1} \sqrt{(x+2)/5}\]

OpenStudy (anonymous):

How to derivative?

OpenStudy (anonymous):

are there 2 equations or one?

OpenStudy (anonymous):

one

OpenStudy (anonymous):

you have to derivate sqrt( (3−x)(2+x)) and derivate 5 arcsins sqrt((x+2)/5) qerivate of sqrt derivate of [(3-x)(2+x)]^1/2 but of arcsin I dont remember the table of derivates look on the table for derivate of arcsin sorry